A student is measuring the average initial rate over the first 10 seconds of the reaction. They collect the following data:

t = 0 s   [A] = 0.131 mol L-1

t = 10 s  [A] = 0.097 mol L-1

What is the initial rate for this reaction?


1
Expert's answer
2021-09-29T05:00:54-0400

Initial rate of the reaction is the instantaneous rate at the start of the reaction; t=0. The initial rate is equal to the negative of the slope of the curve of reactant concentration vs time at t=0.

Initial rate of the reaction "= -\frac{[A_t] -[A_0]}{t-t_0}"

"= -\frac{0.097-0.131}{10-0} \\

= \frac{0.034}{10} \\

= 3.4 \times 10^{-3}"